# Modern approaches to building effective interpretable models of the property market using machine learning

### v1

@@ 5 Modeling
From the correlation matrix (Figure 10) we can see that correlations between the dependent variable and pricing variables for the segment “Flat” are significantly smaller than for the segment “Land parcels”. There is no explicit dependence on the object’s location. Nevertheless, we see that socially important factors such as “Number of shopping malls (within a radius of 1 km)” and “Number of cafes (within a radius of 1 km)” have higher correlations with the per square meter price than other factors, but still insufficient. However, we start with an effort to create a simple and interpretable model – linear regression. This step is necessary according to methodological rules of mass valuation of real estate (Ministry for Economic Development of Russia, 2018). Before modeling we did data preprocessing – we selected the factors (features) included in the model and made data standardization (Konayeva et al., 2024). The standardization is useful in the case of considerably different values of factors in order to reduce them to one scale. As we mentioned earlier the selection of the variables can be done in various ways according to theoretic assumptions and/or assumptions of area under the study. In this paper we used a machine algorithm of selection – SelectKBest (library “sklearn”). This method estimates the dependence of factors with a dependent variable using statistic tests and selects the best factors according to preferred criteria. We used the F-test as a statistical test for estimating the linear dependence between factors.
@@
We selected 8 the most significant pricing variables using SelectKBest – “Area of flat”, “Distance to the water objects (straight-line distance)”, “Distance to the seacoast (straight-line distance)”, “Distance to the closest railway station (straight-line distance)”, “Distance to the area of negatively influencing factors (straight-line distance)”, “Number of cafes (within a radius of 1 km)”, “Distance to the town hall (distance by the roads)”, “Wall material: large-panel”. We created the linear regression model using these factors and taking as the dependent variable the per square meter price without any more transformations. The results of the model are shown in Figure 19.
@@
From Figure 19 we can see that the coefficients of variables are significant (); in general the regression is significant (); the coefficients of variables have right sign according to economic substance; model residuals don’t have autocorrelation (); the value of Condition Number is 2,96, it means that factors don’t contain multicollinearity.
@@
Also, important conditions for using a linear regression are expected value of residuals must be equal to zero (in our case we got the value 3,61e-10) and the dispersion of residuals must be constant (Figure 20) without any trend.
@@
But the residuals of our model are not normally distributed () (Figure 21), which means that estimates got by the ordinary least squares are not optimal – they don’t have minimal dispersion in the class of linear unbiased estimators. Also in this model we have relatively low values of (), which means that the explanatory power of the model is poor.
@@
Figure 22 (a) shows the comparison of predicted and real values of per square meter price. In an ideal case, the cluster of points is expected to be an ellipse stretched out along line y=x, which means that our predictive values should approach real values with the least dispersion. But from Figure 22 (a) we can see that in most cases our predictions overestimate the low values and underestimate the high values.
@@
We summarized the results of the linear regression model in Table 4. We can use it for comparison with other models.
train dataset | test dataset | train dataset | test dataset | Average approximation error (train dataset) | Average approximation error (test dataset) |
0,319 | 0,339 | 17 890,32 | 17 947,46 | 11,7% | 11,9% |
@@
Sometimes the data can include non-linear dependencies that make a linear regression ineffective. In such a situation, we can try other algorithms that work better with non-linear dependencies in data. The most popular method is the “Decision tree” (Díaz-Ramírez et al., 2023). This algorithm forms the hierarchic tree structure that consists of rules “If … then …”. The rules are generated on the base of the training dataset in the process of training. The rules of the algorithm are created on natural language and they are simple for understanding. This is an advantage of the algorithm. But also method has significant shortage – it is prone to overfitting. It means that generated rules will fit only the training dataset, this tree will have bad results on other datasets. The extension of this algorithm called “Random Forest” (Breiman et al., 2017) rectifies this shortcoming. This method creates a set of decision trees. The advantage of it is the high accuracy of prediction and also this algorithm tends to work well with hyperparameters by default. Compared to the “Decision tree” the “Random Forest” is much less prone to overfitting. The shortage of this method is the complicated interpretation of modeling results due to the large number of generated rules.
@@
Yet one algorithm that predicts with good accuracy and possesses better interpretability than “Random Forest” is “RuleFit” (Molnar, 2022; Masís, 2021). The underlying concept of the method is a combination of linear regression and predicted models that are based on decision trees. This model was developed in 2005 and was positioned as a sparse linear regression model (a class of models that after the training can have coefficients with zero weights). This algorithm detects the effects of interaction between features in the form of decision rules. RuleFit automatically generates these functions (rules) from decision trees. Using RuleFit we can choose any algorithm that generates the set of decision trees, for example, random forest or gradient boosting. Generated rules are used in the linear regression model as individual variables. The base model of the algorithm is LASSO (Least Absolute Shrinkage and Selection Operator also known as L1 regularization) (Tibshirani, 1996), because it allows the selection of the most significant rules from the set of rules for including them in the final model. LASSO is similar to ordinary linear regression but there is a supplementary component of regularization in the optimization functional of the model. This component is the penalty for the complexity of the model. Mathematically we can write this as:
@@
where - Residual Sum of Squares; - regularization parameter; - coefficients if models. The goal of minimization is reaching a compromise between errors of regression and the dimension of space of features. Therefore RuleFit is a LASSO regression that takes as input both individual factors and the rules created by the random forest algorithm. These rules describe the interplay of variables. The distinctive characteristic of LASSO is going to zero of some factors as a result of minimization. This procedure determines the most informative factors. The tendency of some factors to disappear is controlled with a regularization parameter that is fitted using cross-validation (procedure of empiric estimation of the generalization ability of the algorithm). More information about adjusting the regularization parameter can be found here (Friedman et al., 2010). Due to the fact that the algorithm automatically selects the most significant factors the additional checking of selected factors is not required. Unlike the Ordinary Least Squares (OLS) method, the LASSO (and consequently RuleFit) does not have any prerequisites. We can estimate the effectiveness of the model by using common evaluation metrics such as (Mean Squared Error), (Mean Absolute Error), (Average approximation error), and (Coefficient of determination).
@@
For building the RuleFit model we used both binary and continuous features. As continuous features, we used the factors from the correlation matrix (Figure 10) where correlation with the dependent variable is 0,15 and more. Factors satisfying this are:
@@
As binary features we used the factors describing the storey of the flat (“the first”, “the last”); the wall material of the building (“bricky”, “large-panel”, “monolithic”, “small-element concrete blocks”; at the same time the wall materials such as “made of timber” and “not determined” were excluded from consideration because they are nonrepresentational and don’t have economic substance). Thus we used 17 features for modeling. We created all kinds of combinations of these factors using RuleFit (we create models with the number of factors from 8 to 12). We created 86632 () models in total. For a selection of the best model we used values of and ; and also we checked the signs of variables according to their economic substance. In order to keep the interpretability of the model we applied a limit on the number of rules as a hyperparameter of the model – 50 rules. For a generation of rules we used the Random Forest algorithm. As a result of LASSO the part of rules were evaluated as insignificant and excluded, and only significant were kept. The final model contains 22 rules and 12 features. They are shown in the Table 6. Due to its volume, we present it in the A.
@@
Unfortunately, RuleFit like any other model has its drawbacks. First, as one can see in Table 6, RuleFit sometimes can keep strange rules that are very similar to each other. This is a drawback of the generation of decision trees by the algorithm and it complicates the interpretation of the model. Second, due to the stochastic nature of the tree’s generation process, RuleFit each time will generate new rules. This problem can be solved by setting the state of the random number generator. Additionally the same randomness will lead to a different set of final rules after changing the training data for the model. Thus the RuleFit is a badly reproduced model and because of it, we evaluate the final results of the model on the whole set of data without splitting it on test and train. But in the process of study we did such splitting and made crossvalidation which gives the average to be 0,58. We present the result of the RuleFit model in Table 5.
Average approximation error | ||
0,6 | 13 518,28 | 8,8% |
From Table 5 we can see that the coefficient of determination is equal to 0,6 which is acceptable for our purposes. Values of MAE and average approximation error of RuleFit are lower than for linear regression.
@@ Appendix A The resulting model for RuleFit
1 | linear | 6713,23 | |
2 | linear | -438,15 | |
3 | linear | -3,14 | |
4 | linear | 6482,54 | |
5 | linear | -2,83 | |
6 | linear | 7435,15 | |
7 | linear | 2,02 | |
8 | linear | 76,72 | |
9 | linear | -0,24 | |
10 | linear | -2771,67 | |
11 | linear | -387,96 | |
12 | linear | -0,08 | |
13 | rule | -26328,13 | |
14 | rule | 29815,42 | |
15 | rule | -6277,76 | |
16 | rule | 17615,75 | |
17 | rule | 13697,61 | |
18 | rule | -6846,8 | |
19 | rule | 8955,08 | |
20 | rule | -4897,04 | |
21 | rule | -4758,11 | |
22 | rule | -7064,57 | |
23 | rule | -3785,21 | |
24 | rule | 4401,15 | |
25 | rule | 6457,31 | |
26 | rule | -3693,24 | |
27 | rule | 2686,18 | |
28 | rule | 5532,75 | |
29 | rule | 4444,61 | |
30 | rule | -898,74 | |
31 | rule | 718,14 | |
32 | rule | -387,44 | |
33 | rule | 399,47 | |
34 | rule | 16,4 |

### v2

@@ 5 Modeling
From the correlation matrix (Figure 10) we can see that correlations between the dependent variable and the pricing variables for the segment “Flats” are significantly smaller than for the segment “Land parcels”. There is no explicit dependence on the object’ location. Nevertheless, we see that socially important factors such as “Number of shopping malls (within a radius of 1 km)” and “Number of cafes (within a radius of 1 km)” have higher correlations with the per square meter price than other factors, although still insufficient. However, we start with an effort to create a simple and interpretable model – linear regression. This step is necessary according to the methodological rules of mass valuation of real estate (Ministry for Economic Development of Russia, 2018). Before modeling we prepared the data by selecting the the factors (features) included in the model and performing the data standardization (Konayeva et al., 2024). The standardization is useful in the case of considerably different values of factors in order to reduce them to one scale. As we mentioned earlier the selection of the variables can be done in various ways according to theoretical assumptions and/or specific aspects of studied area. In this paper we used a machine algorithm of selection – SelectKBest (library “SciKit-learn”). This method estimates the influence of the factors on the dependent variable using statistic tests and selects the best factors according to preferred criteria. We used the -test as a statistical test to estimate the linear dependence between the factors and the per square meter price.
@@
We selected 8 the most significant pricing variables using SelectKBest – “Area of flat”, “Distance to the water objects (straight-line distance)”, “Distance to the seacoast (straight-line distance)”, “Distance to the closest railway station (straight-line distance)”, “Distance to the area of negatively influencing factors (straight-line distance)”, “Number of cafes (within a radius of 1 km)”, “Distance to the town hall (distance by the roads)”, “Wall material: large-panel”. We created the linear regression model using these factors and taking as the dependent variable the per square meter price without any more transformations. The results of the model are shown in Table 5.
@@
From Table 5 we can see that the coefficients of variables are significant (-value 0.05); in general the regression is significant (-value(-statistic) 0.01); the coefficients of variables have right sign according to economic substance; model residuals do not have autocorrelation (1.5 Durbin-Watson 2.5); the value of Condition Number (Gujarati and Porter, 2009) is 2.96, it means that factors do not contain multicollinearity.
@@
Also, important conditions for using a linear regression are the expected value of residuals must be equal to zero (in our case we got the value 3.61e-10) and the dispersion of residuals must be constant (Figure 18) without any trend.
@@
But the residuals of our model are not normally distributed (-value(JB) 0.05) (Figure 19), which means that estimates got by the ordinary least squares are not optimal – they do not have minimal dispersion in the class of linear unbiased estimators. Also in this model we have relatively low values of (), which means that the explanatory power of the model is poor.
@@
Figure 20 (a) shows the comparison of predicted and real values of per square meter price. In an ideal case, the cluster of points is expected to be an ellipse stretched out along line y=x, which means that our predicted values should approach real values with the least dispersion. But from Figure 20 (a) we can see that in most cases our predictions overestimate the low values and underestimate the high values.
@@
We summarized the results of the linear regression model in Table 6. We can use it for comparison with other models on the base of the following metrics: , mean absolute error ()
and average approximation error ()
train dataset | test dataset | train dataset | test dataset | train dataset | test dataset |
0.319 | 0.339 | 17 890.32 | 17 947.46 | 11.7% | 11.9% |
@@
Sometimes the data can include non-linear dependencies that make a linear regression ineffective. In such a situation, we can try other algorithms that work better with non-linear dependencies in the data. The most popular method is the “Decision tree” (Díaz-Ramírez et al., 2023). This algorithm forms the hierarchic tree structure that consists of rules “If … then …”. The rules are generated on the base of the training dataset in the process of training. The rules of the algorithm are created on natural language and they are simple for understanding. This is an advantage of the algorithm. But this method also has a significant shortage – it is prone to overfitting. It means that generated rules will fit only the training dataset, this tree will have bad results on other datasets. The extension of this algorithm called “Random Forest” (Breiman et al., 2017) rectifies this shortcoming. This method creates a set of decision trees. The advantage of it is the high accuracy of prediction and also this algorithm tends to work well with hyperparameters by default. Compared to the “Decision tree”, the “Random Forest” is much less prone to overfitting. The shortage of this method is the complicated interpretation of modeling results due to the large number of generated rules.
@@
Yet one algorithm that predicts with good accuracy and possesses better interpretability than “Random Forest” is “RuleFit” (Molnar, 2022; Masís, 2021). The underlying concept of the method is a combination of linear regression and predicted models that are based on decision trees. This model was developed by Friedman and Popescu (2008) and was positioned as a sparse linear regression model (a class of models that after the training can have coefficients with zero weights). This algorithm detects the effects of interaction between features in the form of decision rules. RuleFit automatically generates these functions (rules) from decision trees. Using RuleFit we can choose any algorithm that generates the set of decision trees, for example, random forest or gradient boosting. Generated rules are used in the linear regression model as individual variables. The base model of the algorithm is LASSO (Least Absolute Shrinkage and Selection Operator also known as L1 regularization) (Tibshirani, 1996), because it allows the selection of the most significant rules from the set of rules for including them in the final model. LASSO is similar to ordinary linear regression but there is a supplementary component of regularization in the optimization functional of the model. This component is the penalty for the complexity of the model. Mathematically we can write this as:
@@
where - Residual Sum of Squares; - regularization parameter; - coefficients if models. The goal of minimization is reaching a compromise between errors of regression and the dimension of space of features. Therefore RuleFit is a LASSO regression that takes as input both individual factors and the rules created by the random forest algorithm. These rules describe the interplay of variables. The distinctive characteristic of LASSO is forcing to zero the weight of some factors as a result of minimization. This procedure determines the most informative factors. The tendency of some factors to disappear is controlled with a regularization parameter that is fitted using cross-validation (procedure of empiric estimation of the generalization ability of the algorithm). More information about adjusting the regularization parameter can be found in Friedman et al. (2010). Due to the fact that the algorithm automatically selects the most significant factors the additional checking of selected factors is not required. Unlike the Ordinary Least Squares (OLS) method, the LASSO (and consequently RuleFit) does not have any prerequisites. We can estimate the effectiveness of the model by using common evaluation metrics such as , , , and (Mean Squared Error):
@@
For building the RuleFit model we used both binary and continuous features. As continuous features, we used the factors from the correlation matrix (Figure 10) where correlation with the dependent variable is 0.15 and more. Factors satisfying this are:
@@
As binary features we used the factors describing the storey of the flat (“the first”, “the last”); the wall material of the building (“bricky”, “large-panel”, “monolithic”, “small-element concrete blocks”; at the same time the wall materials such as “made of timber” and “not determined” were excluded from consideration because they are nonrepresentational and do not have economic substance). Thus we used 17 features for modeling. We created all kinds of combinations of these factors using RuleFit (we create models with the number of factors from 8 to 12). We created 86632 () models in total. For a selection of the best model we used values of and ; and also we checked the signs of variables according to their economic substance. In order to keep the interpretability of the model we applied a limit on the number of rules as a hyperparameter of the model – 50 rules. For generation of rules we used the Random Forest algorithm. As a result of LASSO, part of rules were evaluated as insignificant and excluded, and only the significant were kept. The final model contains 22 rules and 12 features. They are shown in the Table 8. Due to its volume, we present it in the A.
@@
Unfortunately, RuleFit like any other model has its drawbacks. First, as one can see in Table 8, RuleFit sometimes can keep strange rules that are very similar to each other. This is a drawback of the generation of decision trees by the algorithm and it complicates the interpretation of the model. Second, due to the stochastic nature of the tree’s generation process, RuleFit each time will generate new rules. This problem can be solved by setting the state of the random number generator. Additionally the same randomness will lead to a different set of final rules after changing the training data for the model. Thus the RuleFit is a badly reproduced model and because of it, we evaluate the final results of the model on the whole set of data without splitting it on test and train. But in the process of study we did such splitting and made crossvalidation which gives the average to be 0.58. We present the result of the RuleFit model in Table 7.
0.6 | 13 518.28 | 8.8% |
From Table 7 we can see that the coefficient of determination is equal to 0.6 which is acceptable for our purposes. Values of MAE and average approximation error of RuleFit are lower than for linear regression.
@@ 6 Conclusion
Acknowledgments
Authors thank Artyom S. Tanashkin for making the valuable comments and suggestions.
Maps are drawn using OpenStreetMap (OpenStreetMap contributors, 2017) and Python library Folium. For data analysis and modeling we used the following Python libraries (listed in the alphabet order): GeoPy, Matplotlib (Hunter, 2007), NetworkX, NumPy (Harris et al., 2020), OSMnx, pandas (Wes McKinney, 2010), RuleFit (RuleFit contributors, ), SciKit-learn, SciKit GStat (Mälicke, 2021), SciPy, seaborn (Waskom, 2021), statsmodels.
@@ Appendix A The resulting model for RuleFit
: Distance to the closest business activities center (straight-line distance)
@@
1 | linear | 6713.23 | |
2 | linear | -438.15 | |
3 | linear | -3.14 | |
4 | linear | 6482.54 | |
5 | linear | -2.83 | |
6 | linear | 7435.15 | |
7 | linear | 2.02 | |
8 | linear | 76.72 | |
9 | linear | -0.24 | |
10 | linear | -2771.67 | |
11 | linear | -387.96 | |
12 | linear | -0.08 | |
13 | rule | -26328.13 | |
14 | rule | 29815.42 | |
15 | rule | -6277.76 | |
16 | rule | 17615.75 | |
17 | rule | 13697.61 | |
18 | rule | -6846.8 | |
19 | rule | 8955.08 | |
20 | rule | -4897.04 | |
21 | rule | -4758.11 | |
22 | rule | -7064.57 | |
23 | rule | -3785.21 | |
24 | rule | 4401.15 | |
25 | rule | 6457.31 | |
26 | rule | -3693.24 | |
27 | rule | 2686.18 | |
28 | rule | 5532.75 | |
29 | rule | 4444.61 | |
30 | rule | -898.74 | |
31 | rule | 718.14 | |
32 | rule | -387.44 | |
33 | rule | 399.47 | |
34 | rule | 16.4 |